three moles of an ideL monatomic gas are at a temprature of 314 K. Then 2163 of heat is added to the gas, and 720 J of work is done on it. what is the final temperature of the gas?
Answer #1

48 mondatomic gas 3 DR SUS IMRAT brom Isk law of theimclyn mics su=Q-W gives a = 21635 - 7205 su = 2163+720 = 28833 2 1DT = 2